Edgar, Inc. has a materials quantity standard of 4 pounds per finished unit and a material price standard of S8.50 per pound. 228,000 pounds of materials were purchased and used in producing 56,000 units. Edgar, Inc.s direct material quantity variance is a. $34,000 U b. $34,000 F. c. $57,000 U. d. $57,000 F. Seven years ago, Snap Company purchased a parcel of land adjacent to its factory for $15,000. Over the next few years, Snap spent $8,000 demolishing an existing building and making other improvements such that the site was suitable for construction. Snap believes if they were to list the land for sale it would fetch a sales price of approximately $32,000. The county assessor has appraised the land at $28,000 for real estate tax purposes. A neighboring business (anxious to expand) has recently offered Snap $40,000 for the parcel. Assuming Snap does not accept the offer to sell, what are the sunk costs related to this transaction? a. $15,000 b. $23,000 c. $28,000 d. $32,000
Answer #1

Solution 1:

Standard quantity of material for actual production = 56000*4 = 224000 pound

Material quantity variance = (SQ - AQ) * SP = (224000 - 228000) * $8.50 = $34,000 U

Hence option a is correct.

Solution 2:

Sunk cost = Cost of land + cost of demolishing = $15,000 + $8,000 = $23,000

Hence option b is correct.